A woman invested 4 times as much at 3% as she did at 2%. The total amount of interest she earns in 1 year from both accounts is $1,120. How much did she invest at each rate?
Answer by ikleyn(52786)

.
A woman invested 4 times as much at 3% as she did at 2%.
The total amount of interest she earns in 1 year from both accounts is $1,120.
How much did she invest at each rate?
~~~~~~~~~~~~~~~~~

She invested x dollars at 2% and 4x dollars at 3%.


Write the total annual interest equation

    0.02x + 0.03*(4x) = 1120  dollars.


Simplify and find x

    0.02x + 0.12x = 1120

         0.14x = 1120

             x = 1120/0.14 = 8000.


ANSWER.  $8000 invested at 2% and $32000 invested at 3%

Solved.